OpenStudy (anonymous):

The supplement of an angle is 3.5 times as large as the angle. Find the angle.

OpenStudy (anonymous):

140 and 40

OpenStudy (anonymous):

How do I set it up AND work it out?

OpenStudy (anonymous):

ok first i ddid 3.5x + x = 180

OpenStudy (anonymous):

folowwing me so far?

OpenStudy (anonymous):

That's what I had! I got the first part right at least, haha. Next?

OpenStudy (anonymous):

ok then factor out the x so that you have x(3.5+1)=180

OpenStudy (anonymous):

then its 3.5+1=4.5 so then you have 4.5x=180

OpenStudy (anonymous):

then 180/4.5 = x

OpenStudy (anonymous):

you get 40 and you are looking for the larger angle so you just subtract 40 from 180

OpenStudy (anonymous):

dont forget medals XD

OpenStudy (anonymous):

Well... hold on. I'm not used to working it out that way.

OpenStudy (anonymous):

ok np how are you used to it becuase i think i remmber what you are talking about

OpenStudy (anonymous):

How was I able to distribute the x? The answer is correct, I'm just confused. I think they way I was taught was to cancel out the 3.5... or something.

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

Oops comma splice, too.

OpenStudy (anonymous):

oh yea that stuff....they like to think that way...you can just use algebra i ignored my teachers and did it this way ummm illl try to think of how they were trying to teach it to me give me a sec

OpenStudy (anonymous):

Well no, it's fine. I like this way better. It's easier!

OpenStudy (anonymous):

thats what i htought >.> makes you actually think of the process...let me see another one and ill help through it again

OpenStudy (anonymous):

Okay: The complement of an angle is 1.5 times as large as the angle.

OpenStudy (anonymous):

ok so now we have 1.5 this time instead of 3.5 and we have complementary angles instead of supplementary what is your first step?

OpenStudy (anonymous):

So it would be 90-x=1.5x?

OpenStudy (anonymous):

that correct now get it in a way that we can work with it

OpenStudy (anonymous):

Hmm...

OpenStudy (anonymous):

what was the form that i used for the other one? i got all my x's to one side

OpenStudy (anonymous):

Okay... so, 180= x+3.5x?

OpenStudy (anonymous):

no still just 90 but the right side is correct

OpenStudy (anonymous):

OOPS Haha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I have supplementary on the brain, I guess.

OpenStudy (anonymous):

haha np

OpenStudy (anonymous):

So it's 90=x+1.5x?

OpenStudy (anonymous):

right now what is the next step? (i think this is the part that was confusing to you the last time.)

OpenStudy (anonymous):

Well technically there is an understood 1 in front of the first x, so it become 2.5x?

OpenStudy (anonymous):

Saying distribute threw me off last time.

OpenStudy (anonymous):

right we factored an x out

OpenStudy (anonymous):

then what do we have?

OpenStudy (anonymous):

x=36?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now depending on what angle they ask for you might need to subtract that from 90

OpenStudy (anonymous):

Okay, I think it just wants this one. Thank you so much!

OpenStudy (anonymous):

np im happy to help XD